OpenStudy (anonymous):

Solve the following systems of equations using the substitution method. What type of system is it? Name the solution if there is one. x-y=3 x+3y=6

OpenStudy (texaschic101):

x - y = 3 x = y + 3 now sub y + 3 in for x in the other equation x + 3y = 6 y + 3 + 3y = 6 -- subtract 3 from both sides y + 3y = 6 - 3 -- combine like terms 4y = 3 -- divide both sides by 4 y = 3/4 now sub 3/4 in for y in either equation to find x x - y = 3 x - 3/4 = 3 x = 3 + 3/4 x = 12/4 + 3/4 x = 15/4 check... x + 3y = 6 15/4 + 3(3/4) = 6 15/4 + 9/4 = 6 24/4 = 6 6 = 6 (correct) so x = 15/4 and y = 3/4 what type of system ?? I guess it is a two variable system of linear equations.

OpenStudy (anonymous):

i think they are asking if it is constant inconsistent or dependent system? I'm not sure?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

sine they are both 6=6 i would guess consistent system?

OpenStudy (anonymous):

@texaschic101